6.3 web assign

What is the size of the payments that must be deposited at the beginning of each 6-month period in an account that pays 7.2%, compounded semiannually, so that the account will have a future value of $150,000 at the end of 18 years? (Round your answer to the nearest cent.)
Answer #1

\(R=\frac{150000}{\left[\frac{\left(1+\frac{0.072}{2}\right)^{2 \cdot 18}-1}{\left(\frac{0.072}{2}\right)}\right]\left(1+\frac{0.072}{2}\right)}\)

answered by: coolstu